OpenStudy (jenniferjuice):

Does anyone undertsand Solving Quadratic Equations with Complex Solutions?!?!?!!?!?!?!

OpenStudy (kirbykirby):

Just proceed the same way as with non-complex solutions. Apply the quadratic formula in the same way. But when you get a negative square root, like \(\sqrt{-6}\), then this becomes \(i\sqrt{6}\) since \(\sqrt{-1}=i\)

OpenStudy (jenniferjuice):